40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 28 04:57:00 2011, kiwamu, Update, LSC, POX11 demod board broken Screen_shot_2011-10-28_at_2.26.02.png
    Reply  Thu Nov 3 18:41:36 2011, kiwamu, Update, LSC, POX11 demod board : haven't been modified yet 
       Reply  Thu Nov 3 19:58:18 2011, kiwamu, Update, LSC, POX11 demod board : modification done POX11demod.png
          Reply  Fri Nov 4 00:13:24 2011, Koji, Update, LSC, XARM locked with POX11 POX11.pdf
Message ID: 5804     Entry time: Fri Nov 4 00:13:24 2011     In reply to: 5802
Author: Koji 
Type: Update 
Category: LSC 
Subject: XARM locked with POX11 

XARM lock was achieved by POX11_I

Summary:

- The  whitening gains of POX11_I and Q are 42dB so that POX11_I have the same amplitude as AS55_I

- The demod phase of POX11 was adjusted to eliminate the PDH signal from the Q phase. The phase is  -100.5deg.

- In order to lock the XARM with POX11_I_ERR, I had to increase the trigger threshold from 0.1 to 0.2 as the arm was
kicked with the threshold of 0.1.


Method

- Lock the X arm with AS55_I at the XARM configuration.

- Adjusted POX11 demod phase so that POX11_Q is minimized.

- POX I&Q whitening gains were adjusted. When they are 42dB, POX11_I_ERR and AS55_I_ERR have almost the same signal amplitude.
(In reality, POX11_I_ERR has +1dB larger amplitude.)

- Adjusted POX11 demod phase again with better precision.

- Measured transfer function between AS55_I_ERR and POX11_I_ERR. As the sign was opposite, the demod phase was -180deg flipped.

- Tried to lock the arm with POX11_I_ERR. It did not acquire the lock. The arm looked kicked by the servo.

- Increased the trigger threshold from 0.1 to 0.2. Now the arm is locked with POX11_I_ERR.

Attachment 1: POX11.pdf  53 kB  Uploaded Fri Nov 4 01:25:50 2011  | Hide | Hide all
POX11.pdf
ELOG V3.1.3-